# Modeling Spatio-Temporal Transport: From Rigid Advection to Realistic Dynamics

### v3

@@ 3.2 Simulated Examples of a Cyclone
Before discussing the simulated examples, we illustrate how to simulate from the two models introduced in Section 3.1. First, the Distributed FF consists in a convex combination of $n_{v}$ realizations of FFs, each with their specified velocity and wight. Thus, to simulate a Distributed FF one just simulates each FF $n_{v}$ times with Algorithm 1. Moreover, the simulation of an Evolving FF can also be carried out with a modification of Algorithm 1. Specifically, when the space random field $X(\mathbf{s})$ is moved pointwisely in time, instead of shifting it according to a constant velocity, it is moved according to a velocity field $\mathbf{v}(\mathbf{s},t)$. Being it a continuous function of space and time, $\mathbf{v}(\mathbf{s},t)$ is discretized over the space and time grids. For practical implementation, we suggest discretizing $\mathbf{v}(\mathbf{s},t)$ on the same space-time grid as the simulation. When it comes to establishing how large grid $\mathcal{G}^{*}$, should be, one can adopt $\text{v}_{\text{max}}=\max\{\lVert\text{v}_{1}\rVert_{\infty},\lVert\text{v}_{2}\rVert_{\infty}\}$ for varying velocity $\mathbf{v}(\mathbf{s},t)=(\text{v}_{1}(\mathbf{s},t),\text{v}_{2}(\mathbf{s},t))^{T}$, where we determine $\lVert\text{v}_{i}\rVert_{\infty}=\inf\{V\geq 0:|\text{v}_{i}|\leq V\text{ for a.e. }(\mathbf{s},t)\}$, $i=1,2$. We include both algorithms in the supplementary materials.
@@
$\theta_{1}^{t},\dots,\theta_{n_{\theta}}^{t}\stackrel{{\scriptstyle iid}}{{\sim}}f_{R}(\theta;\boldsymbol{\alpha}_{R}),$ (14)
@@
$\mathbf{v}^{t}\sim f_{T}(\mathbf{v};\boldsymbol{\alpha}_{T}).$ (15)
Equation (14) denotes the $n_{\theta}$ angles that define the random rotations of the field, where $f_{R}$ is a probability density function (PDF). Equation (15) describes the velocity $\mathbf{v}^{t}$ with which the entire field is advected at time $t$.
@@
$Z(\mathbf{s},t)$ $=\sum_{i=1}^{n_{\theta}}p(\theta^{t}_{i})X^{\theta^{t}_{i}}_{S}(\mathbf{s}-\mathbf{v}^{t}t).$
Here, $X^{\theta^{t}_{i}}_{S}$ denotes the base random field $X_{S}$ after being rotated by the angle $\theta^{t}_{i}$ around a fixed center of rotation, $C$. The term $(\mathbf{s}-\mathbf{v}^{t}t)$ indicates that this rotated field is evaluated at a coordinate shifted by the velocity vector, effectively modeling advection. The weights $p(\theta^{t}_{i})$ are constructed to form a convex combination, ensuring that $p(\theta^{t}_{i})\geq 0$ and $\sum_{i=1}^{n_{\theta}}p(\theta^{t}_{i})=1$. They are derived from the PDF, $f_{R}$, evaluated at each sampled angle. Specifically, the weight for each rotation is calculated by normalizing the PDF value at that angle:
$p(\theta^{t}_{i})=\frac{f_{R}(\theta^{t}_{i};\boldsymbol{\alpha}_{R})}{\sum_{j=1}^{n_{\theta}}f_{R}(\theta_{j}^{t};\boldsymbol{\alpha}_{R})}.$
This approach ensures that rotations corresponding to angles with higher probability density are given more weight in the final mixture.
To simulate the example in Figure 1, we set $n_{\theta}=10$. For each $t\in\{0,\dots,7\}$, the angles are drawn from a Wrapped Normal distribution, $\theta^{t}_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{WN}(0,5)$, and the velocity is drawn from a multivariate Normal distribution, $\mathbf{v}^{t}\sim\mathcal{N}(\boldsymbol{\mu},\mathbf{I})$, with mean $\boldsymbol{\mu}=(5,0)^{T}$ and the identity as covaraince matrix. The rotation is performed around a fixed center $C=(50,65)$. The x marks in the plot indicate the position of this rotation center relative to the advected field, given by $C^{t}=C+\mathbf{v}^{t}t$. For this simulation, the initial field at $t=0$, $Z(\boldsymbol{s},0)=X(\boldsymbol{s})$, is a fixed, deterministic image rather than a realization of a random field. We deliberately choose a simple shape, a faded circle on a constant background, to provide a clear visual benchmark. This approach allows the model’s intrinsic dynamics to be observed directly and clearly.
@@
$\mathbf{v}(\mathbf{s},t)=\left(\frac{R(\mathbf{s})\cdot\text{cos}(t)}{t+1},\frac{R(\mathbf{s})\cdot\text{sin}(t)}{t+1}\right)^{T},$ (16)
@@
Figure 2: Evolving FF simulation of a cyclone (top row) and the corresponding velocity filed $\mathbf{v}(\mathbf{s},t)$ at every time point (bottom row).
@@ 3.3 Comparison with Numerical Solvers
In this section we show how our proposed methods can simulate solutions of some classes of Partial Differential Equations (PDEs). A comparison with numerical PDE solvers is particularly appropriate, as both our method and these solvers are fundamentally constructed upon a specified advection velocity. This shared foundation provides a direct and rigorous basis for validation. We consider simulations on a spatial grid $\mathcal{G}=[0,100]\times[0,100]$ and time grid $t\in\{0,1,2,3\}$. We use package ReacTran (reactran) to compute the solution of the PDEs. Moreover, to compare the simulated space-time random field $Z(\mathbf{s},t)$ and the solution of the numerical solver $Z_{\text{PDE}}(\mathbf{s},t)$, we use Root Mean Square Error (RMSE)
@@
As mentioned in Section 2.1, the FF corresponds to the solution of the pure transport PDE (2). Thus, the most natural first comparison we make is between our simulated FF and the solution of a transport equation. We simulate a FF with $\mathbf{v}=(20,10)^{\top}$, and we solve (2) with a no-flux boundary condition. The results are shown in Figure 3. We notice that our proposed simulation method closely resembles the numerical solution of the PDE. However, the latter shows numerical diffusion, so that the concentration of the image gets slightly dissipated with time (see for instance LeVeque_2002). This dissipation is quantified in the RMSE plot. the error is zero at $t=0$, as both methods are initialized with the identical field, but grows over time due to the numerical diffusion in the PDE solver. By construction, our suggested method described in Algorithm 1 does not suffer from this issue.
@@
We now turn to the comparison of a Distributed FF with the solution of a transport-diffusion PDE. We show how the superposition of multiple FFs, each with a distinct advection velocity, gives rise to both a mean advective transport and an effective diffusive term. We begin by considering that the field is a superposition of multiple components, each transported by a distinct advection velocity. Let this set of component velocities be denoted by $\{\mathbf{v}_{i}\}_{i=1}^{n_{v}}$, with each velocity occurring with a corresponding probability $\{p_{i}\}_{i=1}^{n_{v}}$. The mean velocity of the system, $\overline{\mathbf{v}}$, is the weighted average of these component velocities, namely $\overline{\mathbf{v}}=\sum_{i=1}^{n_{v}}p_{i}\mathbf{v}_{i}$. The velocity deviations from this mean are then defined as $\Delta\mathbf{v}_{i}=\mathbf{v}_{i}-\overline{\mathbf{v}}$. Conceptually, each component velocities $\mathbf{v}_{i}$ can be thought of as arising from a physical transformation $\mathbf{M}_{i}$ acting on a base velocity $\mathbf{v}$. Similarly to christakos2017spatiotemporal, treating the velocity as the sum of a mean and the fluctuations around is an advantageous framework, as we shall see next.
@@
The left-hand side of (17) represents the pure advection of the field $Z$ at the mean velocity $\overline{\mathbf{v}}$. The term on the right is the divergence of the “fluctuating flux”, which arises from the velocity deviations. The equation in its current form is not closed, as the flux depends on the individual components $Z_{i}$. To create a macroscopic equation for $Z$, we must impose further assumptions.
@@
This derivation provides a formal link between our Distributed FF model and the advection-diffusion PDE.
The identification of the diffusion tensor $\mathbf{C}$ with the covariance of the velocity distribution (Taylor1954) has a physical interpretation. Specifically, diffusion is a macroscopic process of spatial spreading. In the Distributed FF model, the microscopic cause of this spreading is the fact that different additive components of the field are being transported at different velocities. The velocity covariance matrix, namely
$\mathbf{C}_{\mathbf{v}}=\sum_{i=1}^{n_{v}}p_{i}\,\Delta\mathbf{v}_{i}\Delta\mathbf{v}_{i}^{\top}=\sum_{i=1}^{n_{v}}p_{i}\,(\mathbf{v}_{i}-\overline{\mathbf{v}})(\mathbf{v}_{i}-\overline{\mathbf{v}})^{\top},$
@@
is the precise mathematical tool that quantifies the magnitude and orientation of this velocity spread. Therefore, the diffusion tensor $\mathbf{C}$ in the macroscopic PDE is the direct physical manifestation of the velocity covariance, and we can assume that $\mathbf{C}$ is proportional to $\mathbf{C}_{\mathbf{v}}$. This principle allows us to construct a Distributed FF that mimics a desired advection-diffusion process by defining a velocity distribution with the appropriate mean and covariance.
@@
To validate our theoretical framework, we compare our Distributed FF model against a numerical solution of the advection-diffusion equation. The PDE was solved on the same spatial and time grids as in the previous FF example, with a velocity of $\overline{\mathbf{v}}=(20,10)^{T}$ and a diffusion tensor of $\mathbf{C}=\begin{pmatrix}5&0\\ 0&1.5\end{pmatrix}$. The Distributed FF model was constructed by drawing $\{\mathbf{v}_{i}\}_{i=1}^{n_{v}}$ from a multivariate normal distribution whose parameters were set to match the PDE, i.e. $\mathbf{v}_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(\overline{\mathbf{v}},\mathbf{C})$. The weights used are $p_{i}=1/n_{v}$, for each $i=1,\dots,n_{v}$. In Figure 4 we compare single Distributed FF realization with $n_{v}=50$ and the PDE numerical solution, showing qualitative agreement between the two. To quantify the model’s convergence and performance, we carry out a simulation study of $500$ iterations using different numbers of components, $n_{v}=10,50,200$. The right panel of Figure 4 quantifies the model’s performance against the PDE solution. The results clearly indicate that as $n_{v}$ increases, both the median RMSE and its variance decrease across all time points, demonstrating that the Distributed FF model converges to the true solution as more components are included. The shape of the RMSE curves also provides insight into the process. As established in the pure transport case, the error is zero at $t=0$ because both methods are initialized with the identical field. The RMSE then rises as the approximation error accumulates, particularly during the initial phase where the concentration gradients are steepest. Subsequently, the curve plateaus because the diffusion process itself smooths the field over time. As the gradients lessen, the rate at which new error is generated decreases, causing the total error to stabilize.

### v4

@@ 3.2 Simulated Examples of a Cyclone
Before discussing the simulated examples, we illustrate how to simulate from the two models introduced in Section 3.1. First, the Distributed FF consists in a convex combination of $n_{v}$ realizations of FFs, each with their specified velocity and wight. Thus, to simulate a Distributed FF one just simulates each FF $n_{v}$ times with Algorithm 1. Moreover, the simulation of an Evolving FF can also be carried out with a modification of Algorithm 1. Specifically, when the space random field $X_{S}(\mathbf{s})$ is moved pointwisely in time, instead of shifting it according to a constant velocity, it is moved according to a velocity field $\mathbf{v}(\mathbf{s},t)$. As a continuous function of space and time, $\mathbf{v}(\mathbf{s},t)$ is discretized over the space and time grids. For practical implementation, we suggest discretizing $\mathbf{v}(\mathbf{s},t)$ on the same space-time grid as the simulation. When it comes to establishing how large grid $\mathcal{G}^{*}$, should be, one can adopt $\text{v}_{\text{max}}=\max\{\lVert\text{v}_{1}\rVert_{\infty},\lVert\text{v}_{2}\rVert_{\infty}\}$ for varying velocity $\mathbf{v}(\mathbf{s},t)=(\text{v}_{1}(\mathbf{s},t),\text{v}_{2}(\mathbf{s},t))^{\top}$, where we determine $\lVert\text{v}_{i}\rVert_{\infty}=\inf\{V\geq 0:|\text{v}_{i}|\leq V\text{ for a.e. }(\mathbf{s},t)\}$, $i=1,2$. We include both algorithms in the supplementary materials.
@@
$\theta_{1,t},\dots,\theta_{n_{\theta},t}\stackrel{{\scriptstyle iid}}{{\sim}}f_{R}(\theta;\boldsymbol{\alpha}_{R}),$ (14)
@@
$\mathbf{v}_{t}\sim f_{T}(\mathbf{v};\boldsymbol{\alpha}_{T}).$ (15)
Equation (14) denotes the $n_{\theta}$ angles that define the random rotations of the field, where $f_{R}$ is a probability density function (PDF). Equation (15) describes the velocity $\mathbf{v}_{t}$ with which the entire field is advected at time $t$.
@@
$Z(\mathbf{s},t)$ $=\sum_{i=1}^{n_{\theta}}p(\theta_{i,t})X^{\theta_{i,t}}_{S}(\mathbf{s}-\mathbf{v}_{t}t).$ (16)
Here, $X^{\theta_{i,t}}_{S}$ denotes the base random field $X_{S}$ after being rotated by the angle $\theta_{i,t}$ around a fixed center of rotation, $C$. The term $(\mathbf{s}-\mathbf{v}_{t}t)$ indicates that this rotated field is evaluated at a coordinate shifted by the velocity vector, effectively modeling advection. The weights $p(\theta_{i,t})$ are constructed to form a convex combination, ensuring that $p(\theta_{i,t})\geq 0$ and $\sum_{i=1}^{n_{\theta}}p(\theta_{i,t})=1$. They are derived from the PDF, $f_{R}$, evaluated at each sampled angle. Specifically, the weight for each rotation is calculated by normalizing the PDF value at that angle:
$p(\theta_{i,t})=\frac{f_{R}(\theta_{i,t};\boldsymbol{\alpha}_{R})}{\sum_{j=1}^{n_{\theta}}f_{R}(\theta_{j,t};\boldsymbol{\alpha}_{R})}.$
This approach ensures that rotations corresponding to angles with higher probability density are given more weight in the final mixture. Notice that the Distributed FF with time-varying transformations admits a natural interpretation as an expected field. Specifically, the discrete sum (16) can be viewed as a Monte Carlo approximation to the expectation with respect to the rotation angle and velocity, i.e., $\int_{\mathbb{R}^{2}}\int_{-\pi}^{\pi}X_{S}^{\theta}(\mathbf{s}-\mathbf{v}t)\,f_{R}(\theta;\boldsymbol{\alpha}_{R})\,f_{T}(\mathbf{v};\boldsymbol{\alpha}_{T})\,d\theta\,d\mathbf{v}$. This connects directly to the framework of schlather2010some, who considered the expected covariance structure arising from a Gaussian random velocity. Our formulation extends this by incorporating both random rotations and translations, with the discrete representation allowing for non-Gaussian distributions and improving in accuracy as $n_{\theta}$ grows.
To simulate the example in Figure 1, we set $n_{\theta}=10$. For each $t\in\{0,\dots,7\}$, the angles are drawn from a Wrapped Normal distribution, $\theta_{i,t}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{WN}(0,5)$, and the velocity is drawn from a multivariate Normal distribution, $\mathbf{v}_{t}\sim\mathcal{N}(\boldsymbol{\mu},\mathbf{I})$, with mean $\boldsymbol{\mu}=(5,0)^{\top}$ and the identity as covaraince matrix. The rotation is performed around a fixed center $C=(50,65)$. The x marks in the plot indicate the position of this rotation center relative to the advected field, given by $C_{t}=C+\mathbf{v}_{t}t$. For this simulation, the initial field at $t=0$, $Z(\boldsymbol{s},0)=X(\boldsymbol{s})$, is a fixed, deterministic image rather than a realization of a random field. We deliberately choose a simple shape, a faded circle on a constant background, to provide a clear visual benchmark. This approach allows the model’s intrinsic dynamics to be observed directly and clearly.
@@
$\mathbf{v}(\mathbf{s},t)=\left(\frac{R(\mathbf{s})\cdot\text{cos}(t)}{t+1},\frac{R(\mathbf{s})\cdot\text{sin}(t)}{t+1}\right)^{\top},$ (17)
@@
Figure 2: Evolving FF simulation of a cyclone (top row) and the corresponding velocity field
$\mathbf{v}(\mathbf{s},t)$ at every time point (bottom row).
@@ 3.3 Comparison with Numerical Solvers
We now compare our simulation methods with numerical solutions of the corresponding PDEs. A comparison with numerical PDE solvers is particularly appropriate, as both our method and these solvers are fundamentally constructed upon a specified advection velocity. While PDE solvers do not constitute ground truth, they provide a deterministic reference that targets the same dynamics from a different computational angle. Since the FF satisfies the transport equation (2), and the Distributed FF satisfies an advection-diffusion equation under appropriate closure assumptions as we show later in this section, the comparison serves to verify that our stochastic constructions recover the expected behavior.
We consider simulations on a spatial grid $\mathcal{G}=[0,100]\times[0,100]$ and time grid $t\in\{0,1,2,3\}$. We use package ReacTran (reactran) to compute the solution of the PDEs. Moreover, to compare the simulated space-time random field $Z(\mathbf{s},t)$ and the solution of the numerical solver $Z_{\text{PDE}}(\mathbf{s},t)$, we use Root Mean Square Error (RMSE)
@@
As mentioned in Section 2.1, the FF corresponds to the solution of the pure transport PDE (2). Thus, the most natural first comparison we make is between our simulated FF and the solution of a transport equation. We simulate a FF with $\mathbf{v}=(20,10)^{\top}$, and we solve (2) with a no-flux boundary condition. The results are shown in Figure 3. Our proposed simulation method closely resembles the numerical solution of the PDE. However, the latter exhibits numerical diffusion, a well-known artifact of finite-difference schemes for hyperbolic equations (LeVeque_2002), causing the concentration of the image to dissipate slightly over time. This dissipation is quantified in the RMSE plot. The error is zero at $t=0$, as both methods are initialized with the identical field, but grows over time as numerical diffusion accumulates in the PDE solver. By construction, our method samples the advected field directly and does not suffer from this artifact, illustrating a practical advantage for simulating pure transport.
@@
We now turn to the comparison of a Distributed FF with the solution of a transport-diffusion PDE. We show how the superposition of multiple FFs, each with a distinct advection velocity, gives rise to both a mean advective transport and an effective diffusive term. We begin by considering that the field is a superposition of multiple components, each transported by a distinct advection velocity $\mathbf{v}_{i}$ with probability $p_{i}$, as introduced in Definition 3.1. Recall from Section 3.1 the mean velocity $\overline{\mathbf{v}}=\sum_{i=1}^{n_{v}}p_{i}\mathbf{v}_{i}$ and velocity covariance $\mathbf{C}_{\mathbf{v}}=\sum_{i=1}^{n_{v}}p_{i}(\mathbf{v}_{i}-\overline{\mathbf{v}})(\mathbf{v}_{i}-\overline{\mathbf{v}})^{\top}$. We define the velocity deviations $\Delta\mathbf{v}_{i}=\mathbf{v}_{i}-\overline{\mathbf{v}}$. Similarly to christakos2017spatiotemporal, treating the velocity as the sum of a mean and fluctuations is an advantageous framework, as we shall see next.
@@
The left-hand side of (18) represents the pure advection of the field $Z(\mathbf{s},t)$ at the mean velocity $\overline{\mathbf{v}}$. The term on the right is the divergence of the “fluctuating flux”, which arises from the velocity deviations. The equation in its current form is not closed, as the flux depends on the individual components $Z_{i}(\mathbf{s},t)$. To create a macroscopic equation for $Z(\mathbf{s},t)$, we must impose further assumptions.
@@
This derivation provides a formal link between our Distributed FF model and the advection-diffusion PDE. The advection-diffusion equation derived here can be compared to the spatio-temporal SPDE framework presented by carlotto2024. Their model includes a random forcing term on the right-hand side that introduces stochastic variability through the product of temporal white noise and spatial noise. Our Distributed FF formulation differs by having zero on the right-hand side, representing deterministic transport-diffusion dynamics where randomness originates entirely from the distribution of velocities rather than from external forcing. These represent two distinct mechanisms for generating spatio-temporal variability, where our approach models uncertainty in the advective flow itself, while their framework adds fluctuations to otherwise deterministic dynamics. In physical systems, both random advection by varying currents and residual turbulent noise may be present, suggesting these formulations address complementary aspects of spatio-temporal phenomena.
@@
The identification of the diffusion tensor $\mathbf{C}$ with the covariance of the velocity distribution (Taylor1954) has a physical interpretation. Specifically, diffusion is a macroscopic process of spatial spreading. In the Distributed FF model, the microscopic cause of this spreading is the fact that different additive components of the field are being transported at different velocities. The velocity covariance matrix, namely $\mathbf{C}_{\mathbf{v}}$, is the precise mathematical tool that quantifies the magnitude and orientation of this velocity spread. Therefore, the diffusion tensor $\mathbf{C}$ in the macroscopic PDE is the direct physical manifestation of the velocity covariance, and we can assume that $\mathbf{C}$ is proportional to $\mathbf{C}_{\mathbf{v}}$. This principle allows us to construct a Distributed FF that mimics a desired advection-diffusion process by defining a velocity distribution with the appropriate mean and covariance.
To validate our theoretical framework, we compare our Distributed FF model against a numerical solution of the advection-diffusion equation. The PDE was solved on the same spatial and time grids as in the previous FF example, with a velocity of $\overline{\mathbf{v}}=(20,10)^{\top}$ and a diffusion tensor of $\mathbf{C}=\begin{pmatrix}5&0\\ 0&1.5\end{pmatrix}$. The Distributed FF model was constructed by drawing $\{\mathbf{v}_{i}\}_{i=1}^{n_{v}}$ from a multivariate normal distribution whose parameters were set to match the PDE, i.e. $\mathbf{v}_{i}\stackrel{{\scriptstyle iid}}{{\sim}}\mathcal{N}(\overline{\mathbf{v}},\mathbf{C})$. The weights used are $p_{i}=1/n_{v}$, for each $i=1,\dots,n_{v}$. In Figure 4 we compare a single Distributed FF realization with $n_{v}=50$ against the PDE numerical solution, showing qualitative agreement between the two. This agreement is notable because the two methods approach the same advection-diffusion dynamics through fundamentally different mechanisms: the PDE solver discretizes the Laplacian operator directly, while the Distributed FF constructs diffusive behavior from the superposition of fields advected at different velocities.
@@
To quantify the model’s convergence and performance, we carry out a simulation study of $500$ iterations using different numbers of components, $n_{v}=10,50,200$. The right panel of Figure 4 quantifies the model’s performance against the PDE solution. The results clearly indicate that as $n_{v}$ increases, both the median RMSE and its variance decrease across all time points, confirming our theoretical derivation and demonstrating that the approximation error is controllable. The shape of the RMSE curves also provides insight into the process. As established in the pure transport case, the error is zero at $t=0$ because both methods are initialized with the identical field. The RMSE then rises as the approximation error accumulates, particularly during the initial phase where the concentration gradients are steepest. Subsequently, the curve plateaus because the diffusion process itself smooths the field over time. As the gradients lessen, the rate at which new error is generated decreases, causing the total error to stabilize.
@@
3.4 Application to Precipitation Advection: Hurricane Florence
To illustrate the practical relevance of the proposed framework, we apply the FF and Evolving FF models to reproduce the rainfall evolution of during Hurricane Florence, a slow-moving tropical cyclone that affected the Carolinas in mid-September 2018. Making landfall near Wrightsville Beach, North Carolina, Florence was characterized by a collapse in environmental steering currents that reduced its forward velocity to approximately 9.3 km/h. This stalling behavior led to persistent rainbands over the region, resulting in catastrophic freshwater flooding and record-breaking rainfall accumulations (stewart2019florence).
We utilize the NCEP Stage IV precipitation product (nelson2016assessment), a multi-sensor analysis combining radar and gauge observations provided by the National Centers for Environmental Prediction, as source of data. This dataset, available via the R package stars (pebesma2023spatial; stars), provides hourly rain accumulations (kg/m2) at approximately 4 km spatial resolution. The analysis focuses on a 3-hour window from 14:00 to 17:00 UTC on September 14th 2018, immediately following landfall.
We define the spatial grid $\mathcal{G}=[1,87]\times[1,118]$ as an $87\times 118$ pixel lattice, covering an area of approximately 350 km $\times$ 470 km centered over southeastern North Carolina. The temporal domain is discretized into 15-minute intervals, with intermediate frames obtained by linear interpolation between consecutive hourly observations. Let $t\in\{0,\dots,12\}$ denote the time steps, where $t=0$ corresponds to the initialization time at 14:00 UTC, and $t=12$ corresponds to the final evaluation time at 17:00 UTC. The velocity field $\mathbf{v}(\mathbf{s},t)$ is estimated via optical flow using a local block-matching algorithm (horn1981; bowler2004). For each location $(\mathbf{s},t)$, the local velocity vector is determined by finding the displacement maximizing the cross-correlation between consecutive frames within a local neighborhood $\mathcal{B}_{\mathbf{s}}$. Specifically, using a block parameter $b=9$ pixels, this neighborhood is defined as $\mathcal{B}_{\mathbf{s}}=\left\{\mathbf{u}\in\mathcal{D}:\|\mathbf{u}-\mathbf{s}\|_{\infty}\leq 4\right\}$. The estimator is given by:
$\widehat{\mathbf{v}}(\mathbf{s},t)=\arg\max_{(\Delta s_{1},\Delta s_{2})}\,\text{Corr}\left(Z(\mathcal{B}_{\mathbf{s}},t),\,Z(\mathcal{B}_{\mathbf{s}}+(\Delta s_{1},\Delta s_{2}),t+1)\right),$
where $(\Delta s_{1},\Delta s_{2})\in\{-4,\ldots,4\}^{2}$ represent candidate spatial displacements in grid units. To avoid spurious matches in regions of uniform intensity, we assign zero velocity where the standard deviation of $Z(\mathcal{B}_{\mathbf{s}},t)$ is below 0.2 kg/m2, and we retain only displacements achieving a correlation above 0.4. The resulting field is smoothed using a Gaussian kernel with $\sigma=2$ pixels.
For the Evolving FF, the spatially-varying field $\widehat{\mathbf{v}}(\mathbf{s},t)$ is used directly at each time step. Notice that, prior to velocity estimation, each frame is smoothed using a Gaussian kernel with $\sigma=1$ pixel to reduce noise. For the standard FF, we compute a single constant velocity by spatially averaging the initial velocity field:
$\widehat{\overline{\mathbf{v}}}=\frac{1}{|\mathcal{G}|}\sum_{\mathbf{s}\in\mathcal{G}}\widehat{\mathbf{v}}(\mathbf{s},t=0).$
Forecasts are generated by advecting the precipitation field according to the estimated velocities, using bilinear interpolation for sub-pixel displacements. All metrics are computed on the interior of the domain, excluding a 15-pixel buffer from the boundaries to avoid artifacts. This results in an evaluation region of size $57\times 88$ pixels. Performance is quantified using RMSE:
$\text{RMSE}(t)=\sqrt{\frac{1}{|\mathcal{G}|}\sum_{\mathbf{s}\in\mathcal{G}}\left(Z(\mathbf{s},t)-\widehat{Z}(\mathbf{s},t)\right)^{2}},$
and the spatial correlation between the observed field $Z(\mathbf{s},t)$ and the simulated field $\widehat{Z}(\mathbf{s},t)$:
$\text{Corr}(t)=\frac{\sum_{\mathbf{s}\in\mathcal{G}}\left(Z(\mathbf{s},t)-\overline{Z}(t)\right)\left(\widehat{Z}(\mathbf{s},t)-\overline{\widehat{Z}}(t)\right)}{\sqrt{\sum_{\mathbf{s}\in\mathcal{G}}\left(Z(\mathbf{s},t)-\overline{Z}(t)\right)^{2}}\,\sqrt{\sum_{\mathbf{s}\in\mathcal{G}}\left(\widehat{Z}(\mathbf{s},t)-\overline{\widehat{Z}}(t)\right)^{2}}},$
where $\overline{Z}(t)=\frac{1}{|\mathcal{G}|}\sum_{\mathbf{s}\in\mathcal{G}}Z(\mathbf{s},t)$ and $\overline{\widehat{Z}}(t)=\frac{1}{|\mathcal{G}|}\sum_{\mathbf{s}\in\mathcal{G}}\widehat{Z}(\mathbf{s},t)$.
The estimated constant velocity for the FF was $\widehat{\overline{\mathbf{v}}}=(-0.21,-0.08)^{\top}$ grid cells per 15-minute step. This corresponds to a translation speed of approximately 3.6 km/h (0.89 grid cells per hour) toward the south-southwest ($202^{\circ}$). This estimate is notably lower than the overall storm motion of 9.3 km/h reported by stewart2019florence, reflecting the fact that the optical flow captures the slow rotation of spiral bands during the stalling phase rather than just the translation of the eye. Over the 3-hour forecast horizon, this yields a net displacement of roughly 11 km, corresponding to less than three grid cells.
The advantage of the Evolving FF is evident in Figure 5, which displays the estimated velocity field $\widehat{\mathbf{v}}(\mathbf{s},t)$. Unlike the single averaged vector of the Frozen FF, the evolving field captures substantial spatial heterogeneity. The vector plot reveals the presence of rotation of the system, and identifies localized regions of rapid advection associated with convective rainbands.
Figure 5: Estimated velocity field $\widehat{\mathbf{v}}(\mathbf{s},t)$ on September 14, 2018, at four consecutive hours between 14:00 UTC and 17:00 UTC. The arrows indicate flow direction, while the color scale represents the velocity magnitude $\mathbf{v}(\mathbf{s},t)||$.
Figure 6 shows the observed and forecast precipitation fields at hourly intervals: $t=0,4,8,12$, corresponding to 14:00, 15:00, 16:00, and 17:00 UTC. While visual differences are subtle at early lead times, divergence becomes apparent by 17:00 UTC. The FF simulation maintains a pattern closely tied to the initialization, whereas the Evolving FF better captures the evolution of precipitation.
Figure 6: Observed random field (first row), simulated FF (second row), and simulated Evolving FF (third row), at four consecutive hours between 14:00 UTC and 17:00 UTC, on September 14, 2018.
Finally, Table 1 summarizes the quantitative comparison. The Evolving FF consistently outperforms the FF. For instance, at $\Delta t=180$ minutes, the Evolving FF achieves a correlation of 0.76 compared to 0.48 for the FF, and reduces RMSE from 8.7 to 6.4 kg/m2.
We conclude that the advantage of the Evolving FF derives from its ability to accommodate local variations in the flow, which a rigid advection model cannot represent.
Table 1: Performance metrics for Hurricane Florence precipitation forecasting on September 14th, 2018. Lead time $\Delta t$ is measured relative to the initialization at 14:00 UTC.
Time $\text{Corr}(t)$ RMSE (kg/m2)
$\Delta t$ UTC $F$ Evolving FF $F$ Evolving FF
0 min 14:00 1.000 1.000 0.00 0.00
60 min 15:00 0.777 0.832 6.03 5.26
120 min 16:00 0.487 0.628 10.78 9.48
180 min 17:00 0.480 0.761 8.72 6.39